A 3 ft x 6.5 ft rectangular footing is to be embedded 2 ft into the ground and will support a single centrally- located column with the following design loads: P- 50 k, M 80 ft-k (acts in long direction only). The underlying soil is a silty sand with c,-0, ф, 31°-Y-123 pcf, and a very deep GWT. Using a factor of safety of 2.5, determine if this design is acceptable for bearing capacity. Note: Use Vesics bearing capacity factors
Answer #1

.P=sok -M=ook-ft - De=2m C=0, 0 = 31° Y= 123 pcf , FS= 2.5 Now ecentricity, e = M = 80 e=16 bt Now ultimate bearing capacity3 Sq, Sr= Shape factor : - Sa= It B Tano Ź B= B = 3 ft L = L-lez 6.5 - 2x1.6 - 3.3m > B .: Sq= 1+ 3 Tan 31 sq = 1.546 SpaYu = 246 x 20.63 x 1.546x1.19 +0.5X123x3 x 25.99x0.636x11 qu= 12386.36 psf ultimate Load , ou Ou= qu x B x L = 12386.36* 3x 3